OpenStudy (anonymous):

Write a balanced nuclear equation for radioactive decay of thorium-230 with the emission of an alpha particle.

OpenStudy (aaronq):

\(\alpha\) particle = 2 protons + 2 neutrons

OpenStudy (anonymous):

So the equation would be \[\frac{ 4 }{ 2 } He = 2p + 2n\] ?

OpenStudy (aaronq):

you start off with thorium

OpenStudy (anonymous):

I know that it's \[\frac{ 230 }{ 90 } Th\] But where do I put it into the equation?

OpenStudy (aaronq):

on the left side because it's a reactant

OpenStudy (anonymous):

So in that case, \[\frac{ 230 }{ 90 } Th + \frac{ 4 }{ 2 } He = 2p + 2n\] ?

OpenStudy (aaronq):

nope, it's decay, the alpha particle is a product

OpenStudy (anonymous):

So I isolate the alpha product, right?

OpenStudy (aaronq):

idk what you mean by "isolate", the \(\alpha\) goes on the right

OpenStudy (anonymous):

just the a? or with the protons and neutrons too?

OpenStudy (anonymous):

\[\frac{ 230 }{ 90 } Th + 2p + 2n = \frac{ 4 }{ 2 }He\]

OpenStudy (aaronq):

the \(\alpha\) is the 2 protons and 2 neutrons

OpenStudy (anonymous):

I thought a particle is the nucleus of a helium atom?

OpenStudy (aaronq):

it is

OpenStudy (aaronq):

what I'm getting at is that the products are an \(alpha\) particle and the result from subtracting 2 protons and 2 neutrons from the thorium atom

OpenStudy (anonymous):

So in that case... \[\frac{ 230 }{ 90 } Th - 2p - 2n = \alpha particle\]

OpenStudy (aaronq):

you don't have to show a subtraction, otherwise the equation isn't balanced \(^{230}_{90}Th \rightarrow ~^4_2He+~^{226}_{88}?\)

OpenStudy (anonymous):

i understand it now, but why isn't there a Th on the right side?

OpenStudy (anonymous):

so im assuming if it just asked for an equation, I would be allowed to show the subtraction?

OpenStudy (aaronq):

we'll that element decayed into another element, you need to identify it by looking at the number of protons (which is the atomic number).

OpenStudy (aaronq):

the problem is that if you show the subtraction then it wont be balanced (unless you don't show the alpha particle on the right side).

OpenStudy (doc.brown):

Think of it backward for a second. Take radium, cram in a hydrogen nucleus. You get raduim + hydrogen nucleus reacts to form thorium. Chemistry is always concerned with products. Chemists love saying "Matter can not be created nor destroyed", who says nor? Anyway, Thorium reacts to form an alpha particle and what? \[^{230}_{90}Th \rightarrow ~^4_2He+~^{226}_{88}?\] as @aaronq asked

OpenStudy (aaronq):

you mean Helium nucleus

OpenStudy (doc.brown):

yes i do